I need help how to solve
1.
-3x+7y=-16
-9x+5y=16

2.
-4x+9y=9
x-3y=-6
3.
5x+y=9
10x-7y=-18
4.
-6x+6y=6
-6x+3y=-12
5.
8x+y=-16
-3x+y=-5

Answer by ewatrrr(24785)

 
Hi,
1.
Once you have found one of the variables, substitute that into one of the EQs to find the other variable
-3x+7y=-16 |multiply thru by -3 to eliminate the x variable. Then, Add the two EQs to find y
-9x+5y=1
2.
-4x+9y=9
x-3y=-6 |multiply thru by 4 to eliminate the x variable. Then, Add the two EQs to find y
3.
5x+y=9 |multiply thru by -2 to eliminate the x variable. Then, Add the two EQs to find y
10x-7y=-18
4.
-6x+6y=6 |multiply thru by -1 to eliminate the x variable. Then, Add the two EQs to find y
-6x+3y=-12
5.
8x+y=-16 |multiply thru by -1 to eliminate the y variable. Then, Add the two EQs to find x
-3x+y=-5